# Pruning One More Token is Enough: Leveraging Latency-Workload Non-Linearities for Vision Transformers on the Edge

### v2

@@ Abstract
This paper investigates how to efficiently deploy vision transformers on edge devices. Recent methods reduce the latency of transformer neural networks by removing or merging tokens, with small accuracy degradation. However, these methods are not designed with edge device deployment in mind: jthey do not leverage information about the latency vs. workload trends to improve efficiency. First, we show the latency-workload size relationship is nonlinear for certain workload sizes. We consider this relationship to create a token pruning schedule. Second, we demonstrate a training-free, token pruning method utilizing this schedule. We show that for single batch inference, other methods increase latency by 18.6-30.3% with respect to baseline, while we can reduce it by 9%. For similar latency (within 5.2%) across devices we achieve 78.6%-84.5% ImageNet1K accuracy, while the state-of-the-art, Token Merging, achieves 45.8%-85.4%.
@@ 1 Introduction
In the past decade, Internet of Things (IoT) and commodity edge devices have become ubiquitous [4, 22, 43]. Edge devices have become sufficiently powerful, and model miniaturization techniques sufficiently capable, that machine learning (ML) models can be deployed to the network edge for various tasks, including computer vision applications [12, 38]. However, state of the art performance on various computer vision tasks are often claimed by large vision transformer (ViT) based neural network [10] architectures. ViT models such as DeiT [39] and DINOv2 [30] are not designed nor miniaturized for edge deployment. Additionally, latency is often of critical importance on edge devices, and only small inputs or workloads can be processed [18].
Prior work has shown that ViT’s have high redundancy that can be exploited for latency reduction benefits [23]. One general approach is to identify and remove parts of the input (tokens) with low information content. This is called token sparsification. Training-free token sparsification methods such as Token Merging (ToMe) [1] have been effective at reducing latency of pre-trained models. Other approaches like DynamicViT [36] can yield better accuracy than training-free methods, but require training on server-grade hardware.
@@
We address three shortcomings of existing techniques. (1) Many existing efficient methods do not consider fine-grained hardware or latency characteristics [1, 19, 42].  Fig. 1 demonstrates the diversity of latency-workload trends across devices and workload sizes. As a result, there is room to optimize efficient methods in a hardware-aware manner by considering this relationship. (2) Some existing efficient methods may require extensive training [47, 36, 5], hindering the deployment of pre-trained models on edge devices. (3) Prior work has investigated hardware-aware methods for CNN’s [46, 45], but there is little work that shows how to handle or leverage latency-workload behavior for ViT’s. Finally, these works lack primary measurements of underlying GPU or kernel behavior. Thus, our work focuses on reducing ViT latency by considering ViT latency-workload relationships, and without requiring training. This paper has the following contributions:
@@
    We introduce the first method using latency-workload relationships for deciding ViT token pruning schedules.
@@
    We develop a novel training-free token pruning mechanism. For similar latency (within 5.2%) across various hardware and workload sizes, we achieve 0.46 to 43.7 percentage points higher accuracy than ToMe, a state-of-the-art method.
@@ 2.1 Model Acceleration of Vision Transformers
ViT [10] architectures such as DINOv2 [30] have achieved state-of-the-art accuracy on multiple computer vision tasks, including image classification and object detection. State-of-the-art models such as DINOv2-G [30] have over 1 billion parameters. It is important to address the efficiency of ViT models when deploying on edge devices. There are many existing model acceleration techniques for ViT’s such as quantization [21], knowledge distillation [44], low-rank factorization [6], and optimized kernels for attention [8]. In general, these techniques either remove redundant information or reduce network layer compute.
One approach related to redundant information removal is token pruning (sparsification) [1, 17, 37, 36]. These methods are easily applied to a variety of ViT architectures [39, 30, 10]. A potential benefit of pruning approaches is that they do not necessarily require any training or fine-tuning [41, 1, 2]. However, some methods [36, 47, 5, 42] may require significant training time, e.g. 100+ epochs, to recover more accuracy and yield latency reductions. Training poses a high barrier to application; training-free methods are more accessible [1, 41].
@@ 2.2 Latency-Workload Relationship
The thesis of token pruning is that by decreasing workload (tokens), we can reduce latency. However, the relationship between latency and workload size can be non-linear as seen in Fig. 1. This relationship can stem from how workloads are dispatched to the GPU by an ML framework, framework overhead [11], and kernel design decisions [27].  Tab. 1 illustrates primary causes of kernel inefficiency.
@@ 3 Instruction Usage Math Pipe Stalls, Instructions
An example of Cause 1 occurs where a kernel grid size is chosen such that a partial wave of computation must be launched on the GPU — this can lead to a phenomenon termed the GPU tail effect [28]. A partially full wave will incur the same latency cost as a full wave, and this can accumulate on a layer-wise basis. Consequently, a minor change in workload size may lead to a disproportionately large change in latency, since the overhead of partial waves accumulates across layers. For example, on the NVIDIA AGX Orin [24], removing one token can decrease latency by up to 33% (Fig. 1.b).
There are many factors at play that affect latency and kernel behavior. Hardware differences, version-to-version ML framework changes, and the use of closed-source backend libraries like cudNN [7] and cuBLAS [26] make it difficult to model or predict neural network latency behavior. To illustrate this difficulty, we show more examples in Sec. 3.1.
Prior work has exploited the tail effect, which is related to Cause 1, to guide pruning methods for convolutional networks [46, 45]. Kernel-based optimization methods such as FlashAttention [8] may attempt to choose kernel launch configurations that maximize occupancy. Quantization techniques address Cause 2 by using data types that require fewer bits than 32-bit floating point for network parameters. Cause 3 can be addressed by choosing low-level operators that might be faster at the cost of precision, or vice-versa [27].
In this work, we address token pruning in the context of ViT models. Previous work frames CNN channel pruning in the context of Cause-1 problems, specifically the GPU tail effect. However, ViT token pruning mechanisms are fundamentally different from previous CNN channel pruning [14] approaches due to architectural differences between CNN’s and ViT’s. We hypothesize and later demonstrate that latency-workload relationships can also be leveraged to make better token pruning decisions for ViT’s.
@@ 3 Token Pruning with Consideration of Latency and Workload Size
In Sec. 2.1 and Sec. 2.2, we discussed the convenience of training-free token pruning methods and how previous work considered latency and workload size relationships for efficiency benefits. Therefore, we set two design goals: (1) require no training or fine-tuning of a pre-trained ViT model; and (2) achieve more and/or better accuracy-latency tradeoffs by pruning tokens according to these relationships. Tab. 2 illustrates qualitative differences between our work and others as a result of these goals.
@@
Table 2: Qualitative differences between our pruning approach and similar techniques. We consider ViT latency-workload relationships in our pruning, making us hardware aware.
As depicted in Fig. 3, our token pruning approach consists of two main parts. First, we determine a pruning schedule for a particular model and device pair — this includes identifying the number of tokens to remove and the layers at which to remove them. Second, we develop a training-free pruning technique to remove non-informative tokens during inference time. In Sec. 3.1 we show how to decide the number of tokens to prune based on the latency-. Next, in Sec. 3.2 we explain our choice of which layers to prune, completing our offline pruning schedule selection. Finally, Sec. 3.3 describes our token pruning mechanism which is used at inference time.
Figure 2: Latency-workload characteristics of attention operators in PyTorch [34]. Flash, MemEfficient, and Math are optimized, while Vanilla is not. Median latency was measured over 100 runs for each token count. All measurements had an IQR $<$ 1us. The two annotated latency changes of MemEff are discussed in Sec. 3.1.
Figure 3: Illustration of our method to decide a pruning schedule (left), and how we prune according to the schedule at inference time (right).
@@ 3.1 Deciding a Number of Tokens to Prune for our Pruning Schedule
ViT workload size. Before explaining our method we discuss the core operator of ViT models, the attention mechanism [40]. ViT models are parameterized (partially) by an embedding dimension $d$, and their inputs have batch size $b$ and a number of tokens $n$. The input size, or workload, for attention is a tensor with shape $b\times n\times d$. Fig. 1 illustrates the workload size vs latency relationships by varying $n$, $d$, and $b$ across a set of devices.
@@
Measuring latency behavior effects vs. predicting them. Inm As previously mentioned, the GPU tail effect is one phenomenon that arises due to suboptimal kernel grid size choice. Prior work has modeled [20] or utilized [46, 45] GPU tail effect behavior with respect to convolutional neural networks CNN’s. Other work has noted the (sometimes drastic) effect of latency overhead from ML frameworks [11].
Yu et al. [45] notes that the latency of CNN’s increases in discrete jumps, and in all other workload size intervals latency remains the same. In our experiments, we find that internal ML framework operator selection can lead to a range of latency behavior. This claim for CNN’s does not hold in our evaluations for ViT’s. Fig. 2 visually illustrates the latency-workload characteristics of various attention operators in PyTorch [32] with $b$=1 and varying $n$ for a DinoV2-G [30] sized attention operation. This was measured on an NVIDIA RTX 3090 Ti GPU. Additionally, we find that the tail effect is not always the primary factor for drastic latency changes, as was described in previous work. We illustrate these findings with measurements of metrics mentioned in Sec. 2.2 and their co-occurrence with latency behavior.
In one example from Fig. 2, the MemEff attention operator features a $\sim$40latency increase from 128 to 130 tokens that is correlated with a $\sim$40increase in pipeline stall or wait time (Cause 3). But, the kernel grid size remained the same, indicating that the tail effect is not the primary cause of this latency increase (Cause 1). A second example is the $\sim$13latency decrease of MemEff from 132 to 134 tokens is correlated with a 125% increase in kernel grid size (Cause 1) and a 34% decrease in pipeline stall time (Cause 2). These effects seem to be downstream of a different kernel being chosen due to internal heuristics. Importantly, we note latency can even decrease as workload increases, due to underlying kernel, hardware, and framework behaviors.
@@
Ranking token importance. Given a model $M$ with a number of input tokens $N$, we frame choosing the number of tokens to prune $R$ as a multi-objective optimization problem between latency gains and accuracy degradation. This is the offline computation for selecting a pruning schedule for $M$ on a target device, as seen in Fig. 3.
First, we measure latency $L(n)$ and accuracy degradation $A(n)$ of $M$ on the target device for each number of tokens to keep $n\in[1,2,...,N]$. Measuring latency and accuracy of $M$ is performed with a grid search across $n$. We note that in some cases this can be expensive — however we find in many cases it is sufficient to perform the grid search with a stride $>1$ to save time. $A(n)$ only needs to be measured once for a particular $M$, and does not need to be done on the target device. It is only necessary to record latency $L(n)$ on the target device.
For accuracy degradation, we need to define $A(n)$ in a way that does not depend on our pruning schedule selection. Furthermore it is useful to underestimate the accuracy of $M$ so our selection algorithm is hesitant to remove too many tokens, which can degrade accuracy significantly. A simple proxy for the accuracy of each $n$ is to apply random token removal after the first layer on $M$ [31]. It is assumed that any token pruning method should be better than random token removal since random token pruning does not consider token information content at all [42]. Furthermore, pruning at the first layer will degrade accuracy more than pruning later in the network [3]. Therefore random token pruning is a suitable heuristic for estimating accuracy degradation.
Figure 4: Example offline computation to decide the number of tokens prune. We find $n=N_{\text{keep}}$ with maximum utility and by extension the number of tokens to prune $R$ per Eq. 3. Configuration: $\alpha$=0.5, $M$=DinoV2-G, and AGX Orin.
@@ 3.2 Pruning Schedule: Deciding Layers at which to Prune
Now we explain the layers to which our pruning mechanism will be applied. This is the final step of the offline computation illustrated in Fig. 3. Our schedule prunes all $R$ tokens at one layer, early in the network. This differs from other methods such as ToMe [1], Top-K [13, 19], and DynamicViT [36] that progressively prune tokens. Our design is based on two observations supported by our evaluation: First, on small workloads, repeated pruning operations themselves can introduce non-trivial latency (Sec. 4.5). Second, latency benefits compound with each layer after pruning occurs — earlier pruning means more layers will benefit (Sec. 4.3).
Compared with progressive pruning methods such as ToMe [1] that prune at every layer, our approach attempts to place the token pruning at or near the first 25% of layers, similar to the first pruning layer of DynamicViT [36]. For DinoV2-G, we prune at the 10th layer, and for DeiT models, we prune at the 4th layer. This allows us to realize latency reductions according to the GPU tail effect for the remaining 30 layers in DinoV2-G and the 8 remaining layers for DeiT-S and DeiT-B. We discuss the effect of pruning location in Sec. 4.3. As an example of the overall process,  Fig. 4 illustrates an example of the plots of $L(n)$, $A(n)$ and the overall utility $U(n)$. With $N=257$, we follow Eq. 3 to obtain $N_{\text{keep}}=91$ and $R=257-91=166$.

### v3

@@ Abstract
This paper investigates how to efficiently deploy vision transformers on edge devices for small workloads. Recent methods reduce the latency of transformer neural networks by removing or merging tokens, with small accuracy degradation. However, these methods are not designed with edge device deployment in mind: they do not leverage information about the latency-workload trends to improve efficiency. We address this shortcoming in our work. First, we identify factors that affect ViT latency-workload relationships. Second, we determine token pruning schedule by leveraging non-linear latency-workload relationships. Third, we demonstrate a training-free, token pruning method utilizing this schedule. We show other methods may increase latency by 2-30%, while we reduce latency by 9-26%. For similar latency (within 5.2% or 7ms) across devices we achieve 78.6%-84.5% ImageNet1K accuracy, while the state-of-the-art, Token Merging, achieves 45.8%-85.4%.
@@ 1 Introduction
In the past decade, Internet of Things (IoT) and commodity edge devices have become ubiquitous [3, 21, 41]. Edge devices have become sufficiently powerful, and model miniaturization techniques sufficiently capable, that machine learning (ML) models can be deployed to the network edge for various tasks, including computer vision applications [11, 36]. However, state of the art performance on various computer vision tasks is often claimed by large vision transformer (ViT) based neural network [9] architectures. ViT models such as DeiT [37] and DINOv2 [29] are not designed nor miniaturized for edge deployment. Additionally, latency (the time required to do a forward pass given a batch of inputs) is often of critical importance on edge devices, and only small inputs or workloads can be processed [17].
Prior work has shown that ViTs have high redundancy that can be exploited for latency reduction benefits [22]. One approach involves identifying and removing low-information tokens; this is called token sparsification. Training-free token sparsification methods such as Token Merging (ToMe) [1] have been effective at reducing latency of pre-trained models. Other approaches like DynamicViT [34] can yield better accuracy than training-free methods, but require training on server-grade hardware.
@@
We address three key shortcomings in existing techniques. (1) Many existing efficient methods do not consider fine-grained hardware or latency characteristics [1, 18, 40].  Fig. 1 demonstrates the diversity of latency-workload trends across devices and workload sizes. As a result, there is room to improve efficient methods in a hardware-aware manner by considering this relationship. (2) Some existing efficient methods may require extensive training [46, 34, 4], hindering the deployment of pre-trained models on edge devices. (3) Prior work has investigated hardware-aware methods for CNNs [45, 44], but there is little work that shows how to handle or leverage latency-workload behavior for ViTs. Finally, these works lack direct measurements of underlying GPU or kernel behavior. Thus, our work focuses on reducing ViT latency by considering ViT latency-workload relationships, and without requiring training. This paper has the following contributions:
@@
    We propose the first method using latency-workload relationships for deciding ViT token pruning schedules.
@@
    We design a novel training-free token pruning mechanism. For similar latency across various hardware and workload sizes, we achieve 0.46 to 43.7 percentage points higher accuracy than ToMe, a state-of-the-art method.
@@ 2.1 Model Acceleration of Vision Transformers
ViT [9] architectures such as DINOv2 [29] have achieved state-of-the-art accuracy on multiple computer vision tasks, including image classification and object detection. State-of-the-art models such as DINOv2-G [29] have over 1 billion parameters. It is important to address the efficiency of ViT models when deploying on edge devices. Numerous techniques exist for accelerating ViT models, including: quantization [20], knowledge distillation [43], low-rank factorization [5], and optimized kernels for attention [7]. In general, these techniques either remove redundant information or reduce network layer compute.
One approach related to redundant information removal is token sparsification [1, 16, 35, 34]. These methods are easily applied to a variety of ViT architectures [37, 29, 9]. One advantage of sparsification methods is that they often do not necessarily require training or fine-tuning [39, 1]. However, some methods [34, 46, 4, 40] may require significant training time, e.g. 100+ epochs, to recover more accuracy and yield latency reductions. Training poses a high barrier to application; training-free methods are more accessible [1, 39].
@@ 2.2 Latency-Workload Relationship
The premise behind token pruning is that reducing the workload (tokens) can decrease latency. However, this relationship can be non-linear, as demonstrated in Fig. 1. This relationship can stem from how workloads are dispatched to the GPU by an ML framework, framework overhead [10], and kernel design decisions [26].  Tab. 1 illustrates primary causes of kernel inefficiency.
@@ 3 Instruction Usage Math Pipe Stalls, Instructions
An example of Cause 1 occurs where a kernel grid size is chosen such that a partial wave of computation must be launched on the GPU — this can lead to a phenomenon termed the GPU tail effect [27]. A partially filled wave incurs the same latency cost as a fully filled one, and this effect can compound across layers. Thus, even minor adjustments in workload size can result in significant latency changes due to the cumulative overhead of partial waves across layers. For example, on the NVIDIA AGX Orin [23], removing one token (97 to 96) can decrease latency by up to 33% (Fig. 1.b).
There are many factors that affect latency. Differences in hardware, changes across ML framework versions, and reliance on proprietary backend libraries like cudNN [6] and cuBLAS [25] complicate the modeling and prediction of neural network latency. To illustrate this difficulty, we show more examples in Sec. 3.1.
Prior work has exploited the tail effect, which is related to Cause 1, to guide pruning methods for convolutional networks [45, 44]. Kernel-based optimization methods such as FlashAttention [7] may attempt to choose kernel launch configurations that maximize occupancy. Quantization addresses Cause 2 by employing data types with fewer bits than 32-bit floating point for network parameters. Cause 3 can be addressed by choosing low-level operators that might be faster at the cost of precision, or vice-versa [26].
In this work, we address token pruning in the context of ViT models. Previous work frames CNN channel pruning in the context of Cause-1 problems, specifically the GPU tail effect. However, ViT token pruning mechanisms are fundamentally different from previous CNN channel pruning [13] approaches due to architectural differences between CNNs and ViTs. We hypothesize and later demonstrate that latency-workload relationships can also be leveraged to make better token pruning decisions for ViTs.
@@ 3 Token Pruning with Consideration of Latency and Workload Size
In Sec. 2.1 and Sec. 2.2, we discussed the advantages of training-free token pruning methods and how previous work considered latency and workload size relationships for efficiency benefits. Therefore, we set two design goals: (1) require no training or fine-tuning of a pre-trained ViT model; and (2) achieve better accuracy-latency tradeoffs by pruning tokens according to these relationships. Tab. 2 illustrates qualitative differences between our work and others as a result of these goals.
@@
Table 2: Qualitative differences between our pruning approach and similar techniques. We consider ViT latency-workload relationships to guide our token pruning.
As depicted in Fig. 3, our token pruning approach consists of two main parts. First, we establish a pruning schedule for specific model-device pairs, which involves determining the number of tokens to prune and the layers where pruning occurs. Second, we devise a training-free technique for pruning non-informative tokens at inference time. In Sec. 3.1 we show how to decide the number of tokens to prune based on the latency-workload relationship. Next, in Sec. 3.2 we explain our choice of which layers to prune, completing our offline pruning schedule selection. Sec. 3.3 describes our token pruning mechanism which is used at inference time. Last, Sec. 3.4 clarifies qualitative differences between our method and existing approaches.
[…]
Figure 3: Illustration of our method to decide a pruning schedule (left) and how we prune according to the schedule at inference time (right), which are discussed in the sectiions shown at the top of the illustration.
@@ 3.1 Deciding a Number of Tokens to Prune for our Pruning Schedule
ViT workload size. Before explaining our method we discuss the core operator of ViT models, the attention mechanism [38]. ViT models are partially parameterized by an embedding dimension $d$, with inputs characterized by a batch size $b$ and a number of tokens $n$. The input size, or workload, for attention is a tensor with shape $b\times n\times d$. Fig. 1 demonstrates how varying $b$, $n$, and $d$ affects the relationship between workload size and latency across various devices.
@@
Measuring latency behavior effects vs. predicting them. As previously mentioned, the GPU tail effect is one phenomenon that arises due to suboptimal kernel grid size choice. Prior work has modeled [19] or utilized [45, 44] GPU tail effect behavior with respect to convolutional neural networks CNNs. Other work has noted the (sometimes drastic) effect of latency overhead from ML frameworks [10].
Yu et al. [44] claims that the latency of CNNs increases in discrete jumps, and in all other workload size intervals latency remains the same. In our experiments, we find that internal ML framework operator selection can lead to a range of latency behavior. This claim for CNNs does not hold in our evaluations for ViTs. Fig. 2 depicts the latency-workload characteristics of various attention operators in PyTorch [31] with $b$=1 and varying $n$ for a DinoV2-G [29] attention operation. This was measured on an NVIDIA RTX 3090 Ti GPU. Additionally, we find that the tail effect is not always the primary factor for drastic latency changes, as was described in previous work. We illustrate these findings with measurements of metrics mentioned in Sec. 2.2 and their co-occurrence with latency behavior.
In one example from Fig. 2, the MemEff attention operator features a $\sim$40latency increase from 128 to 130 tokens that is correlated with a $\sim$40increase in pipeline stall or wait time (Cause 3). But, the kernel grid size remained the same, indicating that the tail effect (Cause 1) was not the primary cause of this latency increase. A second example is the $\sim$13latency decrease of MemEff from 132 to 134 tokens is correlated with a 125% increase in kernel grid size (Cause 1) and a 34% decrease in pipeline stall time (Cause 2). These effects seem to be downstream of a different kernel being chosen due to internal heuristics. Importantly, we note latency can even decrease as workload increases, due to underlying kernel, hardware, and framework behaviors.
@@
Ranking token importance. Given a model $M$ with $N$ input tokens, we define the selection of $R$ tokens to prune as a multi-objective optimization problem, balancing latency gains against accuracy degradation. This is the offline computation for selecting a pruning schedule for $M$ on a target device, as seen in Fig. 3.
First, we measure latency $L(n)$ of $M$ on the target device for each number of tokens to keep $n\in[1,2,...,N]$. Measuring latency and accuracy of $M$ is performed with a grid search across $n$this is demonstrated in Algorithm 1. Running times for all configurations we evaluate over are listed later in Sec. 4.2.
{minted}
[fontsize=]python def measure_latency(model, b, N): """ b refers to batch size, N is the number of tokens ’model’ expects """ #L(n), as a dictionary L = for n in range(1, N): # Latency is independent of random inputs # model.d is the embedding size (Sec. 3.1) x = torch.rand(b,n,model.d) # Benchmark for fixed time latency = bench(model, x) L[n] = latency return L
Algorithm 1 Offline Workload Latency Measurement
{minted}
[fontsize=]python def measure_accuracy(model, dataset, N): """ dataset is the eval split, N is the number of tokens ’model’ expects """ #A(n), as a dictionary A = for n in range(1, N): # Running accuracy for n n_acc = 0.0 for image,label in dataset: # Shape (b, N, model.d) x = model.embed(image) # Shape (b, n, model.d) x = random_prune(x, n) y_pred = model.predict(x) n_acc += sum(y_targ == label) A[n] = n_acc / len(dataset) return A
Algorithm 2 Offline Accuracy Degradation Estimate
Second, we estimate the accuracy $A(n)$ of $M$ after pruning tokens. We need to measure $A(n)$ in a way that does not depend on our pruning schedule selection, however. Furthermore it is useful to underestimate the accuracy of $M$ so our selection algorithm is hesitant to remove too many tokens, which can degrade accuracy significantly. A simple proxy for the accuracy of each $n$ is to apply random token removal after the first layer on $M$ [30], which we refer to as random_prune. Algorithm 2 depicts how to compute $A(n)$. It is assumed that any token pruning method should be better than random token removal since random token pruning does not consider token information content at all [40]. Furthermore, pruning at the first layer will degrade accuracy more than pruning later in the network [2]. Therefore random token pruning is a suitable heuristic for estimating accuracy.
@@
The following definitions meet these criteria:
$\text{U}_{\text{L}}(n)$ $=1-\cfrac{L(n)}{\text{max}\,[\,L(n)\,]}$ (1)
@@ 3.2 Pruning Schedule: Deciding Layers at which to Prune
We explain where in the ViT our pruning mechanism is applied. Our schedule prunes all $R$ tokens at one layer, early in the model. This differs from other methods such as ToMe [1], Top-K [12, 18], and DynamicViT [34] that progressively prune tokens. This choice is based on two observations supported by our evaluation: First, on small workloads, the repeated application of pruning operations can introduce significant latency (Sec. 4.3). Second, latency reductions accumulate with each subsequent layer after pruning; thus, pruning earlier allows more layers to benefit from low latency.
{minted}
[fontsize=]python def vit_forward(model, x, N, R, L): """ x is an image-like input. N is the number of tokens after embedding x. R is the number of tokens to prune, and L is the layer at which to prune (Sec. 3.2). """ # Tokenize input image, x has shape (b, N, d) x = model.embed(x) for idx, layer in enumerate(model.layers): # Standard self-attention x, attn, V = layer(x) if idx == L: scores = rank_tokens(attn, V) x = prune_tokens(x, scores, N, R) return model.head(x)
def rank_tokens(attn, V): """ attn and V are from self-attention. attn has shape (b, h, N, N), where h is the number of attention heads. V has shape (b, h, N, d / h). ’keepdim=True’ means the tensor dimension reduced over is not removed, but is kept with length=1. """ am = max(attn,dim=1).sum(dim=1,keepdim=True) am /= max(am.transpose(-2,-1)) # V metric (ours) vm = max(V, dim=1).sum(dim=-1,keepdim=True) vm = softmax(vm, dim=1) # Tokenwise scores with shape (b, N, 1) return am + vm
def prune_tokens(x, scores, N, R): """ x is a tensor of tokens. We prune such that N-R tokens remain. """ # Sort by score, descending sorted_scores_idx = argsort(scores,dim=1) # Shape (b, N-R-1, d) kept = gather(x, sorted_scores[:,:N-R-1]) # Shape (b, 1, d) inattentive = gather(x, sorted_scores[:,N-R-1:]).mean(dim=1) # Shape (b, N-R, d) return torch.cat([kept, inattentive], dim=1)
Algorithm 3 Inference-Time Pruning Mechanism
We perform pruning after the first 25% of ViT layers, akin to the first pruning layer of DynamicViT [34] — this yields latency reduction for the remaining 75% of layers. We refer to the index of this pruning layer as $L$. Different pruning locations are evaluated in the supplemental work.